# I Introduction

### v1

@@ Abstract
We present the results of an experimental investigation into buoyant rigid spheres rising through concentrated collections of hydrated hydrogel particles. The volume fraction of particles is such that the mechanical properties of the material are intermediate between a very viscous fluid and a soft solid. Despite the non-Newtonian character of these hydrogels, we find that when the surface of the material is free, an immersed buoyant sphere rises with a constant speed. The effects of the motion in the surrounding material are observed to be highly localized around the sphere. When the mass of the sphere is varied, its terminal velocity is observed to depend exponentially on the buoyancy. Qualitatively distinct behavior is found when a lid is present on the surface of the material, and in this case, a sublinear time dependence of the displacement is found. Linear motion of the sphere is accompanied by flow at the surface of the material whereas fluid movement is suppressed when a lid is applied. Other boundary condition variations are consistent with the perspective that constant speed motion is induced by the presence of a free boundary.
@@ I Introduction
The use of a falling-bob viscometer to make estimates of the viscosity of Newtonian fluids is a standard laboratory procedure. It involves applying Stokes law for the motion past a sphere [1] to the terminal speed of rigid particles falling or rising in very viscous fluids or at very small lengthscales. Stokes law applies when where the Reynolds number with , the terminal speed of the sphere of radius and is the kinematic viscosity of the fluid. Typically, the law is valid for many viscous oils and fluids such as glycerine and syrups. The falling bob viscometer technique is nowadays highly refined [2, 3] where effects from the walls of the container [4, 5] or the finite length of the apparatus [6] are taken into account. Moreover, this configuration has been used to investigate the motion of rigid spheres through suspensions of neutrally buoyant spheres over a range of volume fractions [7] and suspensions of fine rods [8]. The method has also been extended to falling cylinders in narrow tubes [9].
@@
Typically, it is assumed that the fluid properties are independent of time and speed for falling-bob viscometers to be used effectively. However, these techniques can also be used to provide deeper insights into the flow behavior of some non-Newtonian fluids, such as viscoelastic materials [10]. A study using a shear thinning fluid found that the distance the sphere travels is [11]. An explanation for this non-linear behaviour is provided by [12] who demonstrate that a build up of stagnant fluid around the sphere leads to an increase in its effective size and hence Stokes drag increases with time. In shear thickening fluids, a similar coupling has been observed between the effective size of the intruder and the viscous response of the material surrounding it for the non-linear slowdown of an intruder [13].
@@
The focus of our current study is on a class of non-Newtonian materials, namely mixtures of hydrated hydrogel particles. Indicative of the ambiguous mechanical nature of these materials, we also refer to them as hydrogel suspensions when we discuss their overall fluid-like properties, or hydrogel packing when we discuss the static properties of the sedimented hydrogel particles. They have many remarkable features, including sensitivity to the frictional properties of the hydrogels, non-Newtonian flow profiles, critical packing fraction behavior in their high flow rate rheology and various forms of creep behavior [14, 15, 16, 17, 18]. Control of the properties of hydrogel suspensions in water can be achieved via the friction of hydrogel particles [14], packing density [19, 20, 16, 17] and even using the valency of added ions [21]. In previous studies, the non-Newtonian fluid properties of dense hydrogel suspensions is paramount. For example, hydrogel suspensions have been shown to behave as Herschel-Bulkley fluids [14] when probed with standard rheological methods under steady state and constant driving rate conditions.
@@ I.1 Constant speed motion
Our primary goal is to use the conceptually simple geometry of a falling-bob viscometer to investigate our counterintuitive observation that under some circumstances a buoyant, rising sphere can travel at a constant speed through samples of hydrated hydrogel suspensions. Results are obtained over a wide variety of sample conditions and sphere sizes and are surprising since hydrogel suspensions are materials with well-established non-Newtonian properties. Indeed, our previous work revealed a creep flow motion in which the speed typically decreases with the square root of time [17, 18], indicating the presence of significant non-Newtonian behavior at these constant stress driving conditions. Hence, observation of constant speed motion motion over a wide parameter range is a new and surprising result.
In our investigations, we systematically vary the boundary condition of the hydrogel particle mixture to find that the difference between the constant speed motion observed in the current work and the nonlinear creep motion observed previously originates in the boundary conditions applied to the material. Despite the rheology of hydrogel suspensions being an intrinsic property of the microscopics of the material, we establish that its global mechanical response can be affected by the boundary conditions, which is highly unusual. Our work suggests that further theoretical work is required to describe the constitutive behaviour of hydrogel particle suspensions in the presence of specific boundary conditions.
@@ I.2 Dependence on material properties and boundary conditions
The dependence on temperature of the observed suspension flow behavior is observed to be significantly greater than the temperature dependence of water alone. This finding indicates that the effective viscosity of the suspension is not dominated by the solvent viscosity alone e.g. via a local mechanism such as a lubrication layers between the particles, as these would scale sinking or rising speeds proportionally with the water viscosity. In fact, particle image velocimetry analyses reveal that the deformation field of the hydrogel suspension is highly localised around the intruder.
@@ II.1 Upper Boundary Conditions
The cylinder is sealed at the top by a tight fitting lid containing a neoprene ’O’ ring. The sealed container helps to minimise any evaporation of the water. There is a cm gap between the upper surface of the sample and the underside of the external lid. The upper boundary of the sample is a free surface for of the experiments and a printed plastic lid is placed on the surface of the sample in the remainder of the experiments. The lid is 3D printed in plastic (PLA), hence it is partially aerated and weighs g. Its bottom surface is smooth since it is formed in contact with the heating element of the 3D printer, and is hence comparatively free of defects. It is placed on the surface of the hydrogel packing and a thin layer of water is squeezed on onto the upper surface. The lid is lightweight and the results of an investigation into the effects of confinement [18] indicate its contribution to the bulk stress on the material is minimal.
@@ II.2 Particle tracking
The motion of the sphere is tracked by capturing images using a computer controlled camera at set intervals ranging from to seconds depending on the buoyancy of the ball. In the case shown in Fig.1 the images are taken at second intervals using a Nikon digital camera. The recorded images are stored on a PC using the software DigiCam and are processed offline using the software ImageJ. The speed is estimated from the resulting distance versus time plots.
@@ II.3 Sample preparation
Sample preparation proceeds as follows. Dry hydrogel powder is slowly mixed with the water which is gently stirred and left to settle for two hours between additions. The particles swell as they absorb water and sink to the bottom of the cylinder as they are slightly more dense than the fluid. The sample is gently stirred between each charge. The hydrogels absorb the water efficiently and form a layer of swollen particles at the bottom of the cylinder with a clear layer of fluid above. Measurements of the depth of the hydrogel layer as a function of the added mass of dry powder is linear as in previous experiments [17]. This finding is consistent with the hydrogel particles being essentially frictionless spheres [25, 14, 26], which settle to a random close packing. The hydrogel particle packing reaches the surface of the water at a dry hydrogel content g/L where is defined as the weight of the dry powder divided by the total volume of water in grams per Litre (g/L). Once the layer has reached the surface, the addition of further quantities of dry hydrogels does not give any detectable change in the layer height and presumably there is a small compression of the particles throughout the layer. The average size of the swollen particles is measured to be mm. The value of used in the majority of the experiments reported here is set to g/L.
@@
The experiments with rising spheres were performed over a period of approximately one year in a laboratory constructed in TM’s home which is a well-insulated property with a thatched roof. This provides a good short-term stable temperature environment. The temperature of the sample is recorded at regular intervals and for the majority of the period it is found to be °C. Over the summer months the temperature varied outside of this range and the effects of this are reported below. However, the majority of the data given below is collected with the temperature within the given range. We will see below that this temperature window gives at most a variation of measured travelling speeds, which is insignificant for the effects considered, given the large range of travelling speeds explored via the variation of buoyancy stress and intruder size.
@@
Experiments were also performed in a laboratory environment at the University of Amsterdam, under standard laboratory temperature control. In these cases a suspension of hydrogel spheres was formed in a square plexiglass box of dimensions cm following our established protocol using dry hydrogel spheres with triply boiled Amsterdam tap water at a density of g/L unless otherwise noted. To initiate the experiments, a rod with a hoop was mainly used to place the sphere at the bottom, and in other cases a small magnet was placed inside the ball. An external magnet was used to hold and release the ping pong ball from the bottom of the container.
@@
In the Amsterdam experiments the software TroublePix was used with a Lumenera Lt225C high speed camera to record a video of the motion at a frame rate of frames/second. Still images were captured using a Nikon D5200 camera with ball tracking imaging performed using MaxTraq and Matlab. The flow field was measured using Particle Image Velocimetry (PIV). In this, the plane of motion of the sphere was illuminated by two vertical green laser sheets. Spatial averaging which co-moves with the intruder was used. Matlab integrated software “PIVlab” was used to analyse the sequences of images.
@@ II.4 Intruder specifications
Ping pong balls are manufactured from Acrylonitrile Butadiene Styrene (ABS) plastic, are mm in diameter and weigh g. Their size, sphericity and weight are strictly controlled by the International Table Tennis Federation [27] and the dimensions of each ball are checked using a vernier gauge. The diameter ratio of the sphere to container where is the diameter of the sphere and is that of the cylinder. Changing the buoyancy of a ping-pong ball is relatively straightforward as water can be injected into the ball using a hypodermic needle, the injection hole sealed and the weight of the ball measured using an electronic balance. The buoyancy force can then be estimated using the difference in mass between the sphere and the displaced water. We assume that the stress that an intruder exerts on the hydrogel suspension is its effective weight divided by its projected cross sectional area. The main set of experiments were performed using ping-pong balls with a hydrogel sample of 6.35 g/L. A second set of experiments were carried out using a hollow 20mm diameter polypropylene ball (Precision plastic ball co.). Injecting water and weighing was again used to change and estimate the buoyancy. The value of was 6.0 g/L. A third set of experiments were performed with hollow plastic balls. The balls again came from the Precision Ball Co., were made from a range of plastics (polyamide, polypropylene, PTFE, polyethylene) and had diameters mm. These hollow plastic balls were weighed using an electronic balance and the the weight was used to compute the buoyancy stress. The value of used was 6.15 g/L for these experiments. Finally, in one experiment, a glass sphere with a diameter of 25mm was used.
@@ III Overall rising intruder dynamics
The indicative yet informative and simple result of an initial set of experiments is shown in Fig. 2
where it can be seen from a single typical experimental data set that the distance a sphere rises in a hydrogel sample is proportional to when there is a free surface at the top of the sample. The inset Fig. 2 shows the linearity of the displacement dynamics also on a double logarithmic scale.
The constant speed motion stands in stark contrast with previous square-root time dependence observations [17, 18]. From the systematic exploration of intruder motion as reported below, we will see that the (end) boundary plays a defining role in the timescale of the motion of the sphere through the sample. We first present a systematic investigation into this constant speed motion of buoyant spheres moving towards a free boundary.
@@ III.1 Time Histories with a Free Surface
As discussed above, the buoyancy of the ping-pong ball can readily be changed by adding small measured quantities of water to its interior. The time histories obtained over a range of buoyancies are shown in Fig. 3 (a). It can be seen that a straight line provides a satisfactory least squares fit to each time history with typically only a 3% root mean squared error, and hence there is a linear dependence of distance on time. A wide range of buoyancies was explored using this approach, from a sphere without added water (Pa) and the total rise time is minutes to one with Pa, where the sphere takes hours to rise through the sample. We show all the behaviour for the experiments with ping-pong balls in a sample with g/L and a free upper surface in Fig. 3 (a). A fixed terminal velocity is found for each value of in the specified sample. This suggests an interpretation of the results using the laws of standard Stokes viscometry might be possible. Of course, the connection cannot be complete since we shall see that the rise velocity depends exponentially on buoyancy; the hydrogel sample has clear non-Newtonian properties. Even so, models of rigid particles in suspensions [28] suggest that an effective viscosity can be ascribed to suspensions of particles in fluids. Following this approach we extract an effective viscosity of the hydrogel medium from the constant intruder velocity data with
@@
where is the effective dynamic viscosity of the fluid, is the radius of the sphere, is the difference in density between the intruder and hydrogel medium respectively, is gravity and is the terminal velocity of the sphere. Strictly speaking, this formula is only valid for a Stokes flow in an unbounded domain. If the fluid medium used here was a Newtonian viscous fluid, wall and end effects would be significant for the geometry used here. However, as was found in [29] and reported below, the hydrogel suspension motion is only significant close to the sphere, so we choose the simplified analysis for the present study. We will show PIV measurements below which also indicate that the flow field is localised around the sphere.
@@
The effective viscosity approach allows us to rescale all the acquired displacement data for one sample onto a master curve that follows . We show the result of this analysis in Fig. 3 (b). An excellent collapse over three orders of magnitude confirms the constant speed dynamics of an intruder rising towards a free surface.
@@
The collapse of the data provides an important message. It is possible that the time dependence observed previously [17, 18] could arise from a gradient in the properties of the sample resulting from concentration variations or the effects of hydrostatic pressure, for example, or other gradients because of the sensitivity of the viscosity to packing density around the random close packing limit [19]. The constant speed motion, instead, highlights that concentrated samples of hydrated hydrogels which do not display spatiotemporal density variation effects can be consistently produced.

### v2

@@ Abstract
We present the results of an experimental investigation into buoyant rigid spheres rising through highly concentrated collections of hydrated hydrogel particles. The volume fraction of particles is such that the mechanical properties of the material are intermediate between a very viscous fluid and a soft solid. Despite the established time dependent, non-Newtonian character of hydrogels, we find that when the surface of the material is free, an immersed buoyant sphere rises with a constant speed. The effects of the motion are observed to be highly localized around the sphere. When the stress exerted on the material is changed by varying the mass of the sphere, its terminal velocity is found to depend exponentially on its buoyancy. Qualitatively distinct behavior is found when a solid lid is placed on the surface of the material. In this case, a seemingly thixotropic, sublinear time-dependent motion is found. It is observed that linear motion of the sphere is accompanied by flow at the surface of the material whereas fluid movement is suppressed when a lid is present. We use these observations to provide a hypothesis which links the exponential stress dependence of the rheology of the material to the effects of the boundary conditions on the kinematics of the intruder.
Contents
@@ I Introduction
We report the results of an experimental investigation into the slow motion of rigid spheres through a soft solid. The soft material comprises a highly concentrated collection of hydrated hydrogels. The particle concentration is such that the swollen particles form an interconnected network of highly elastic, porous material with no visible regions of water in the interior or at the surface of the sample. The buoyancy driven motion of the sphere involves complex interactions determined by viscoelastic and capillary forces so that the material response is a mixture of fluid-like deformation and solid-like resistance. Applications of such behaviours are to be found in soft robotics and soil sciences [1].
@@
By way of analogy we make connections with the falling bob viscometer which is a standard laboratory method for measuring the viscosity of very viscous fluids. The method involves applying Stokes law for the motion past a sphere [2] to the terminal speed of rigid particles falling or rising in very viscous fluids or at very small lengthscales. Stokes law applies when where the Reynolds number with , the terminal speed of the sphere of radius and is the kinematic viscosity of the fluid. In practice, the law is valid for many viscous oils and fluids such as glycerine and syrups. The falling bob viscometer technique is nowadays highly refined [3, 4] where effects from the walls of the container [5, 6] or the finite length of the apparatus [7] are taken into account. Moreover, this configuration has been used to investigate the motion of rigid spheres through suspensions of neutrally buoyant spheres over a range of volume fractions [8, 9] and suspensions of fine rods [10]. The method has also been extended to falling cylinders in narrow tubes [11].
@@
Typically, it is assumed that the fluid properties are independent of time and speed for falling-bob viscometers to be used effectively. Moreover, in the case of suspensions, there is a tacit assumption that the kinematics of the particles in the viscous fluid is a direct metric of the local Newtonian rheology of the material. In this case, the local shear stress is proportional to the local shear rate via the viscosity. However, falling bob viscometers can also be used to provide deeper insights into the flow behavior of non-Newtonian fluids, such as viscoelastic materials [12, 13]. For example, in a study of a shear thinning fluid it is found that the distance the sphere travels is [14]. An explanation for this non-linear behaviour is provided by [15] who postulate that a build up of stagnant fluid around the sphere will lead to an increase in its effective size over time and hence Stokes drag increases accordingly. In shear thickening fluids, a similar coupling has been observed between the effective size of the intruder and the viscous response of the material surrounding it for the non-linear slowdown of an intruder [16], which may lead to oscillatory velocity fluctuations [17]. These examples highlight that for non-Newtonian fluids, the observed intruder kinematics can only be reliably interpreted when a detailed knowledge of the constitutive relationship of the fluid is used in combination with an appropriate model of the boundary conditions.
@@
The focus of our current study is on a class of non-Newtonian materials, namely mixtures of hydrated hydrogel particles. Indicative of the ambiguous mechanical nature of these materials, we also refer to them as hydrogel suspensions when we discuss their overall fluid-like properties, or hydrogel packing when we discuss the static properties of the sedimented hydrogel particles. They have many remarkable features, including sensitivity to the frictional properties of the hydrogels, non-Newtonian flow profiles, critical packing fraction behavior in their high flow rate rheology and various forms of creep behavior [18, 19, 20, 21, 22]. Control of the properties of hydrogel suspensions in water can be achieved via the friction of hydrogel particles [18], packing density [23, 24, 20, 21] and even using the valency of added ions [25]. In previous studies, the non-Newtonian fluid properties of dense hydrogel suspensions is known to be paramount. For example, hydrogel suspensions have been shown to behave as Herschel-Bulkley fluids [18] when probed with standard rheological methods under steady state and constant driving rate conditions.
@@ I.1 Constant speed motion
Our primary goal is to use the conceptually simple geometry of a falling-bob viscometer to investigate our counterintuitive observation that under some circumstances a buoyant, rising sphere can travel at a constant speed through samples of hydrated hydrogel suspensions, despite their apparent thixotropic properties. Results are obtained over a wide variety of sample conditions and sphere sizes and are surprising since hydrogel suspensions are materials with well-estabished non-Newtonian properties. Indeed, our previous work revealed a creep flow motion in which the speed typically decreases with the square root of time [21, 22], indicating the presence of significant time dependent, non-Newtonian behavior at these constant stress driving conditions. Hence, observation of constant speed motion that depends exponentially on the driving stress over a wide parameter range is a novel and surprising result which reinforces the need for establishing a connection with the constitutive equations that govern soft sphere suspensions.
In our investigations, we systematically vary the boundary conditions of the hydrogel particle mixture to find that the difference between the constant speed motion observed in the current work and the nonlinear creep motion observed previously originates in the boundary conditions applied to the material. Despite the rheology of hydrogel suspensions being an intrinsic property of the microscopics of the material, we establish that its local kinematics can be affected by the boundary conditions, which is highly unusual. Our work suggests that further theoretical work is required to describe how the constitutive behaviour of hydrogel particle suspensions is affected by the presence of specific boundary conditions. Below, we offer a hypothesis that connects the exponential stress sensitivity in the rheology to the change in thixotropic behavior under different boundary conditions.
@@ I.2 Dependence on material properties and boundary conditions
The dependence on temperature of the observed suspension flow behavior is observed to be significantly greater than the temperature dependence of water alone. This finding indicates that the effective viscosity of the suspension is not dominated by the solvent viscosity alone e.g. via a local mechanism such as lubrication layers between the particles, as these would scale sinking or rising speeds in proportion with the water viscosity. In fact, particle image velocimetry analyses reveal that the deformation field of the hydrogel suspension is highly localised around the intruder.
@@ II.1 Upper Boundary Conditions
The cylinder is sealed at the top by a tight-fitting lid containing a neoprene ’O’ ring. The sealed container helps to minimise any evaporation of the water. There is a cm gap between the upper surface of the sample and the underside of the external lid. The upper boundary of the sample is a free surface for of the experiments and a printed plastic lid is placed on the surface of the sample in the remainder. The lid is 3D printed in plastic (PLA), hence it is partially aerated and weighs g. Its bottom surface is smooth since it is formed in contact with the heating element of the 3D printer, and is hence comparatively free of defects. It is placed on the surface of the hydrogel packing and a thin layer of water is squeezed on onto the upper surface. The lid is lightweight and the results of an investigation into the effects of confinement [22] indicate its contribution to the bulk stress on the material is minimal.
@@ II.2 Particle tracking
The motion of the sphere is tracked by capturing images using a computer-controlled camera at set intervals ranging from to seconds depending on the buoyancy of the ball. In the case shown in Fig.1 the images are taken at second intervals using a Nikon digital camera. The recorded images are stored on a PC using the software DigiCam and are processed offline using the software ImageJ. The speed is estimated from the resulting distance versus time plots.
@@ II.3 Sample preparation
Sample preparation proceeds as follows. Measured amounts of dry hydrogel powder are slowly mixed with the water which is gently stirred and left to settle for two hours between additions. The particles swell as they absorb water and sink to the bottom of the cylinder as they are slightly more dense than the fluid. The sample is gently stirred between each charge. The hydrogels absorb the water efficiently and form a layer of swollen particles at the bottom of the cylinder with a clear layer of fluid above. Measurements of the depth of the hydrogel layer as a function of the added mass of dry powder is linear as in previous experiments [21]. This finding is consistent with the hydrogel particles being essentially frictionless spheres [29, 18, 30], which settle to a random close packing. The height of hydrogel particle interface reaches the surface of the water at a hydrogel concentration g/L where is the weight of the dry powder divided by the total volume of water in grams per Litre (g/L). Once the layer has reached the surface, the addition of further quantities of dry hydrogels does not give any detectable change in the layer height and presumably there is a small compression of the particles throughout the layer. The average size of the swollen particles is measured to be mm.
The value of used in the majority of the experiments reported here is set to g/L. We choose to use the concentration to set the properties of the sample. As mentioned above, this is a robust control parameter which produces repeatable experimental results. The more common parameter of volume fraction used in studies of suspensions of hard spheres [31] is difficult to estimate with soft materials. Recently, it has been shown that accurate estimates of the volume fraction of hydrogels in suspension [32] can be made only with specialized imaging equipment, such as X-ray facilities.
@@
The majority of the experiments were performed over a period of approximately one year in a laboratory where the measured temperature of the sample is °C. Over some periods the temperature varied outside of this range and the effects of this are reported below. However, the majority of the data given below is collected with the temperature within the given range. We will see below that this temperature window gives at most a variation of measured travelling speeds, which is insignificant for the effects considered, given the large range of travelling speeds explored via the variation of buoyancy stress and intruder size.
@@
Experiments were also performed at the University of Amsterdam, under standard laboratory temperature control. In these cases, a suspension of hydrogel spheres was formed in a square plexiglass box of dimensions cm following our established protocol using dry hydrogel spheres with triply boiled Amsterdam tap water at a concentration of g/L unless otherwise noted. To initiate the experiments, a rod with a hoop was mainly used to place the sphere at the bottom, and in other cases a small magnet was placed inside the ball. An external magnet was used to hold and release the ping pong ball from the bottom of the container.
@@
In the Amsterdam experiments the software TroublePix was used with a Lumenera Lt225C high speed camera to record a video of the motion at a frame rate of frames/second. Still images were captured using a Nikon D5200 camera with ball tracking imaging performed using MaxTraq and Matlab. The flow field was measured using Particle Image Velocimetry (PIV). In this, the plane of motion of the sphere was illuminated by two vertical green laser sheets. Spatial averaging which co-moves with the intruder was used. Matlab integrated software “PIVlab” was employed to analyse the sequences of images.
@@ II.4 Intruder specifications
Ping pong balls are manufactured from Acrylonitrile Butadiene Styrene (ABS) plastic, have a diameter of mm and weigh g. Their size, sphericity and weight are strictly controlled by the International Table Tennis Federation [33] and the dimensions of each ball are checked using a vernier gauge. The diameter ratio of the sphere to container where is the diameter of the sphere and is that of the cylinder. Changing the buoyancy of a ping-pong ball is relatively straightforward as water can be injected into the ball using a hypodermic needle, the injection hole sealed and the weight of the ball measured using an electronic balance. The buoyancy force can then be estimated using the difference in mass between the sphere and the displaced water. We assume that the stress that an intruder exerts on the hydrogel suspension is its effective weight divided by its projected cross sectional area. The main set of experiments are performed using ping-pong balls with a hydrogel sample with g/L. A second set of experiments were carried out using a hollow mm diameter polypropylene ball (Precision plastic ball co.). Injecting water and weighing is again used to change and estimate the buoyancy. The value of is g/L. A third set of experiments are performed with hollow balls made from a range of plastics (polyamide, polypropylene, PTFE, polyethylene) and had diameters mm. These hollow plastic balls were weighed using an electronic balance and the the weight was used to compute the buoyancy stress. The value of used was g/L for these experiments. Finally, in one experiment, a glass sphere with a diameter of mm was used.
@@ III Overall rising intruder dynamics
The results of a typical set of experiments are shown in Fig. 2 for a representative data set. It may be seen that the distance a sphere rises in this hydrogel sample is proportional to when there is a free surface at the top of the sample. The inset Fig. 2 confirms the linearity of the displacement dynamics on a double logarithmic scale. The constant speed motion stands in stark contrast with previous square-root time dependence observations [21, 22]. From the systematic exploration of intruder motion reported below, we deduce that the (end) boundary plays a defining role in the timescale of the motion of the sphere through the sample. We first present a systematic investigation into the constant speed motion of buoyant spheres moving towards a free boundary.
@@ III.1 Time Histories with a Free Surface
As discussed above, the buoyancy of the ping-pong ball can readily be changed by adding small measured quantities of water to its interior. The time histories obtained over a range of buoyancies are shown in Fig. 3 (a). It can be seen that a straight line provides a satisfactory least squares fit to each time history with typically a 3% root mean squared error, and hence there is a linear dependence of distance on time. A wide range of buoyancies was explored using this approach, from a sphere without added water (Pa) and the total rise time is minutes to one with Pa, where the sphere takes hours to rise through the sample. We show all the behaviour for the experiments with ping-pong balls in a sample with g/L and a free upper surface in Fig. 3 (a).
A fixed terminal velocity is found for each value of in the specified sample. This suggests an interpretation of the results using the laws of standard Stokes viscometry might be possible. Of course, the connection cannot be complete since we shall see that the rise velocity depends exponentially on buoyancy; the hydrogel sample has clear non-Newtonian properties. Even so, models of rigid particles in suspensions [12, 34] suggest that an effective viscosity can be ascribed to suspensions of particles in fluids. Following this approach we extract an effective viscosity of the hydrogel medium from the constant intruder velocity data by converting the directly measured terminal velocity of the rising sphere into a value with the units of Pas, using the traditional Stokes variables as the radius of the sphere, is the difference in density between the intruder and hydrogel medium respectively and for gravity:
@@
In this interpretation, is the effective dynamic viscosity of the fluid. Strictly speaking, this formula is only valid for a Stokes flow in an unbounded domain. If the fluid medium used here were a Newtonian viscous fluid, wall and side-wall effects would be significant for the geometry used here. However, as was found in [35] and reported below, motion in the hydrogel suspension is only significant close to the sphere, so we choose to use the simplified analysis for the present study. We will show PIV measurements below, which also indicate that the flow field is localised around the sphere.
@@
The effective viscosity approach enables us to rescale all the displacement data acquired for one sample onto a master curve that follows . We show the result of this analysis in Fig. 3 (b). An excellent collapse over three orders of magnitude confirms the constant speed dynamics of an intruder rising towards a free surface.
@@
The collapse of the data provides an important insight. It is possible that the time dependence observed previously [21, 22] could arise from a gradient in the properties of the sample resulting from concentration variations or the effects of hydrostatic pressure, for example, or other gradients because of the sensitivity of the viscosity to packing density around the random close packing limit [23]. The constant speed motion, instead, highlights that concentrated samples of hydrated hydrogels which do not display spatiotemporal density variation effects can be consistently produced.